OpenStudy (anonymous):

What is the new volume of a rectangular prism that is originally twelve cubic meters and is then resized using a scale factor of 4 on each side length?

OpenStudy (anonymous):

HELP ME PLEASE ITS IMPORTANT!!!

OpenStudy (anonymous):

@bakonloverk @mebs @litchlani Do you guys know this? please help me!

OpenStudy (ybarrap):

Since this is a volume, scaling each side by a factor \(a\), will increase its volume by a factor \(\bf \large a^3\). Take the volume of a cube: $$\large \tt V_{orig}=x\times y\times z\\ \text{Let's scale each side by a factor a:}\\ \large\hspace{10pt}V_{new}=ax\times ay\times az\\ \large\hspace{10pt}V_new=a^3V_{orig} $$ The factor you have is 4. So, how much more volume will your new prism have after scaling?

OpenStudy (ybarrap):

Your rectangular prism has a volume \(depth\times width \times height\), similar to what I have above. Each dimension is similarly scaled by the factor, 4. So multiply the original volume by \(\tt 4^3\).

OpenStudy (anonymous):

Oh my gosh! Thank you so much! @ybarrap

OpenStudy (ybarrap):

yw